OpenStudy (anonymous):

The graph of which of the following will be parallel to the graph of y = 2x – 4?

OpenStudy (anonymous):

A. 2x + y = 4 B. 4x – 2y = –12 C. y = –4x + 8 D. y = 3x – 4

OpenStudy (anonymous):

If we put the answers into y = mx + b can then tell what the slope is for every answer, and only one has the same slope. Parallel lines also all have the same slope. Let me know if you still need more help. A) y = -2x + 4 B) y = 2x + 6 C) y = –4x + 8 D) y = 3x – 4

OpenStudy (anonymous):

yea so whats the anwser im getting confused @S_Stones

OpenStudy (anonymous):

m = the slope, and parallel lines have the same slope. all the answers are now in for form of y = mx +b. so the slope for A) is -2, B) is 2, C) is -4 and D) is 3, we know that we are looking for a slope of 2 (y=2x-4) thus it would be b

OpenStudy (anonymous):

sorry for the late reply, didn't notice the notification.

OpenStudy (anonymous):

do you know how to turn 4x – 2y = –12 into y = 2x + 6 ?

OpenStudy (anonymous):

no @S_Stones

OpenStudy (anonymous):

ok.. so that would be where you are getting lost then, to solve this we first need to put all the answers into a form where we can find the slope. to do they we simply solve for y. so... 4x-2y = -12 -4x = -4x get y by itself ( -4x from both sides) -2y = -4x -12 \-2 = \-2 get y alone and positive (divide each side by -2) y = 2x + 6 we know that it is in Slope-intercept form (y = mx + b) and we know that while in this form m = slope and b = the y intercept. we also know that parallel lines have the SAME slope. thus the answer must have a slope of 2 (y = 2x – 4) hope this helped. if not let me know where you are getting lost.